sweetburger:

Derivation of the 0th order integrated rate law.

Vocaloid:

I just remember that it's rate = k lol

sweetburger:

So an example 0th order integrated rate law can be said as A--> products where \[rate=k[A]^0\] rate can be expressed in this manner \( \frac{d[A]}{dt} \) however we know that this is a rate of disappearance of reactant so its correctly expressed as \( \frac{-d[A]}{dt} \) we can now say \[\frac{ -d[A] }{ dt }=k \] through a simple separation of variables we can say \[\int\limits_{[A]_0}^{[A]_t}d[A]=\int\limits_{0}^{t}-kdt \] \[([A]_t -[A]_0)=-k(t-0)\] \[[A]_t -[A]_0=-kt \] \[[A]_t=-kt+[A]_0\]

sweetburger:

tbh idk why i keep this in my head

Vocaloid:

good work I haven't look at this stuff since freshman chem tbh

sweetburger:

thanks, and as you probably know the idea extends easily to higher order reactions its just a different result of the integration